An electron revolves around a positive charges rod with linear charge density 2×10^4 .Find the kinetic energy of electron if it is independent of radius ?

Grade:12

1 Answers

Eshan
5 years ago
Electric field at a distancerfrom the rod=\dfrac{\lambda}{2\pi\epsilon_0 r}
Hence attractive force acting on the electron =\dfrac{\lambda e}{2\pi\epsilon_0 r}

This force provides the centripetal acceleration to the electron.

\implies \dfrac{\lambda e}{2\pi\epsilon_0 r}=\dfrac{mv^2}{r}
\implies \dfrac{1}{2}mv^2=\dfrac{\lambda e}{4\pi \epsilon_0}